A plane crashes 330 miles from shore. The survivors climb in a raft and drift toward shore at 4 mph. A coast guard ship was sent to find them at the same moment they began to drift. The ship travels at 80 mph.
a) How long until the ship meets the raft?
b) How far from the shore are they?

Answer by Alan3354(69443)

A plane crashes 330 miles from shore. The survivors climb in a raft and drift toward shore at 4 mph. A coast guard ship was sent to find them at the same moment they began to drift. The ship travels at 80 mph.
a) How long until the ship meets the raft?
b) How far from the shore are they?
---------------------
a) The combined speed is 4 mph + 80 mph = 84 mph
330/84 = 55/14 hours = 3 hours & 55.7 mins
-------------
b) The boat travels 80 mph for 55/14 hours
Distance = 80*55/14 miles
= 314.29 miles
BTW, that's a damn fast boat.